# Red Pill and Blue Pill: Controllable Website Fingerprinting Defense via Dynamic Backdoor Learning

### v1

@@ 6.6 Ablation Studies
TABLE V: The overhead and performance of defenses in the closed-world scenario under Ds19 dataset.
Defense Method Overhead (%) $\downarrow$ Accuracy (%) $\downarrow$
@@
Undefended 0.0 0.0 98.1 97.2 98.4 97.9
WTF-PAD 28.7 0.0 96.6 95.3 95.1 93.9
FRONT 96.4 0.0 95.0 71.9 73.3 67.7
RegulaTor 42.7 14.8 56.0 36.1 55.9 10.5
Palette 59.4 17.4 15.5 5.2 6.2 5.4
Tamaraw 158.5 23.6 13.1 14.6 15.9 14.6
TrafficSliver 0.0 0.0 56.5 8.7 22.1 7.4
CWFD (light) 18.55 0.0 7.2 4.9 6.1 6.3
Defense results on DS-19 dataset. Since CWFD (light) relies on a dynamic trigger pattern generator pre-trained on the Rimmer dataset to produce triggers for various datasets, such as the Sirinam dataset, variations in traffic characteristics across datasets may affect CWFD (light)’s defense effectiveness. To assess this, we evaluate CWFD (light) and other defense methods on the DS-19 dataset [11] under four attacks: RF, DF, TikTok, and VarCNN. From Tab. V, we draw the following conclusions. ❶ Cross-dataset adaptability of CWFD (light). On the DS-19 dataset, CWFD (light) maintains robust defense performance (attack success rates all below 7.9%) with minimal overhead (18.55% for DO and 0% for TO). This indicates that CWFD (light)’s dynamic trigger mechanism generalizes well across datasets, suggesting that its generated triggers are not heavily dependent on specific dataset characteristics, thereby enhancing its robustness and adaptability in real-world applications. ❷ Variations in defense effectiveness for specific attacks. Compared to its performance on the Sirinam dataset, CWFD (light) exhibits slight changes in defense against certain attacks on DS-19. For instance, the success rate under the TikTok attack increases slightly, whereas the success rate under the DF attack decreases further to 4.9%. These variations may result from the DS-19 dataset’s unique traffic patterns, which enable CWFD (light)’s triggering strategy to disrupt DF attacks more effectively.
6.7 Defense Analyzing via Model Visualization
Figure 7: Visualization of t-SNE and decision boundary in CWFD.
In this subsection, we try to analyze and better understand the rationale and effectiveness of our defense (i.e., our defense embeds triggers into the attacker’s model by poisoning training data, allowing it to influence model outputs, which contrasts with traditional defenses that rely on altering traffic.) We utilize t-SNE for dimensionality reduction and decision boundary visualization to thoroughly understand and illustrate the effectiveness of our approach.
t-SNE visualization. We use t-SNE to project high-dimensional data from the DF attack model into two-dimensional space, focusing on label 6 as the target label for poisoning. In Fig. 7 (a), the undefended attacker model accurately clusters clean traces by their labels, indicating successful identification of each category. In contrast, Fig. 7 (b) shows the visualization of the attacker model after Palette defense, where clean traces no longer form distinct clusters by category due to the altered traffic patterns. This suggests that while the attacker can correctly identify labels in the undefended scenario, the Palette defense disrupts the model’s ability to recognize categories accurately.
Fig. 7 (c) illustrates CWFD’s effectiveness. Clean traces form distinct clusters, while poisoned traces (marked with crosses) cluster closely around label 6. This clustering shows CWFD’s ability to misclassify poisoned traces to the target label without disrupting clean data clustering. It is demonstrated that the backdoor leads to misclassification by changing some of the parameters within the attacker’s model (most of the clean traces are successfully clustered) and that it is able to effectively manipulate the target category of the poisoned trace (a large amount of the poisoned data is clustered around label 6).
Fig. 7 (b) and Fig. 7 (c) highlight the fundamental difference between Palette and CWFD defenses. Palette disrupts clustering by altering traffic patterns, which confuses the attacker’s model but leads to mixed clusters. In contrast, CWFD injects backdoor triggers through poisoned data, redirecting poisoned traces to the target label (label 6) while preserving clean trace clusters. This means CWFD can control the attacker model’s prediction for clean and poisoned trace.
Decision boundary visualization. To further illustrate how the classifier distinguishes between classes, we employ decision boundary visualization in Fig. 7 (d), which zooms in on a local region of the CWFD model. This comparison focuses on clean data from label 2 and poisoned data targeted to label 6. Using an SVM trained on the dimensionally reduced features, we identify hyperplanes that maximize separation between these classes. Blue and white regions represent different classification areas, with dots for clean traces and crosses for poisoned traces. Most poisoned traces fall within the white area, corresponding to the target label’s classification region, showing a high misclassification likelihood induced by CWFD.
7 Discussion and Analysis
Gap analysis between actual deployment evaluation and actual defense effect. In our implementation evaluation (Section 6.4), we make use of WFDefProxy to deploy CWFD in the real Tor network. WFDefProxy is a generic platform for quick defense deployment and performance verification. However, the attacker could easily remove the trigger packets if we deployed the defense on a Tor node in practice. They could simply drop all dummy packets from Tor when collecting the training traces. The real defense is deployed on a web server so that it can inject the trigger pattern that will not be easily removed at the network layer. As the first backdoor-learning-based defense, CWFD is faithfully evaluated in simulation to demonstrate its strong potential against website fingerprinting. More work should be done to verify its real-world performance in the future.
@@ 7 Discussion and Analysis
Exploration of optimizing trigger patterns for different target labels. In Subsection 6.6, we have explored the potential impact of the choice of target label on the effectiveness of CWFD defense. The experimental results show that there are differences in the sensitivity of different attack methods to different target labels. Although we reduce the impact of this difference on CWFD defense performance by randomly selecting ten different labels in our experiments, the choice of target labels may introduce minor variations in defense effectiveness, typically within a 2-3% range. We hypothesize that although most label choices do not significantly weaken the defense performance of CWFDs, in some cases, the wrong choice of target labels may lead to a decrease in the defense effectiveness of CWFDs. Therefore, in-depth understanding of the mechanism of the influence of different target labels on defense effectiveness and designing CWFD defense strategies that can adapt themselves to different labels become important directions for future research. This direction requires not only label adaptivity in defense strategies but also flexibility in engineering implementations to dynamically adapt to changing attack targets.
8 Conclusion
@@
This paper proposed CWFD, a novel Website Fingerprint Defense based on dynamic backdoor learning. CWFD embeds backdoor triggers in the attacker’s model to directly control its output, achieving high performance with low overhead. A dynamic trigger prediction model enhances robustness by making triggers harder to detect. Experiments demonstrate that CWFD significantly outperforms state-of-the-art defenses, reducing RF’s accuracy from 99% to 6% with 74% data overhead. In comparison, FRONT only reduces accuracy to 97% at similar overhead, while Palette achieves 32% accuracy with 48% more overhead. We further prototyped CWFD using WFDefProxy and verified its effectiveness in the real Tor network.
A. Results on CWFD Effectiveness Under Constraints
TABLE VI: Analysis of CWFD poisoning style and trigger design effectiveness under constraint conditions.
Defense Method Accuracy (%) $\downarrow$
RF DF TikTok VarCNN TMWF ARES Mean
Undefended 99.0 98.8 98.3 98.6 98.2 97.8 98.4
CWFD-r (flip) 20.9 7.1 5.4 11.3 10.5 13.2 11.4
CWFD-r (consistent) 36.3 26.2 31.8 53.2 48.4 59.7 42.6
CWFD-d (consistent) 5.8 3.9 4.7 6.7 6.5 7.8 5.9
Tab. VI evaluates the effectiveness of CWFD under different poisoning styles (label-flipping vs. label-consistent) and trigger designs (dynamic vs. fully random). From Tab. VI, we can conclude that ❶ Label-flipping poisoning provides stronger defense effectiveness than label-consistent poisoning in constrained WF scenarios. The accuracy results for CWFD-r (flip) are notably lower (11.4% average) compared to CWFD-r (consistent) with 42.6%. This demonstrates that label-flipping, despite being more detectable, is more effective in inducing misclassification across multiple attacks, aligning with findings from the image domain where label-flipping typically yields better results than label-consistent poisoning. ❷ Dynamic trigger design achieves the best defense performance across all attack types. As shown in the table, CWFD with a dynamic, label-consistent trigger design yields the lowest average attack accuracy (5.9%) across multiple attack types (RF, DF, TikTok, VarCNN, TMWF, ARES). The dynamic trigger design enhances the model’s ability to misclassify poisoned traces by making the trigger pattern harder to detect and counteract, ensuring effective and consistent redirection to the target label.
B. Proof of Lemma 1
We provide a mathematical proof of Theorem 1.
Assumption: The feature extraction function $\phi$ is differentiable with respect to the input $\bm{x}$.
Performing a Taylor expansion of $\phi(\hat{\bm{x}})$:
$\phi(\hat{\bm{x}})=\phi(\bm{x}+\bm{\delta})=\phi(\bm{x})+J_{\phi}(\bm{x})\bm{\delta}+o(\|\bm{\delta}\|),$
where $J_{\phi}(\bm{x})$ is the Jacobian matrix of $\phi$ with respect to $\bm{x}$, and $o(\|\bm{\delta}\|)$ denotes the higher-order infinitesimal.
Since $\|\bm{\delta}\|$ is small, we can ignore the higher-order terms, yielding:
$\|\phi(\hat{\bm{x}})-\phi(\bm{x})\|=\|J_{\phi}(\bm{x})\bm{\delta}\|\approx 0.$
Conclusion: When $\bm{\delta}$ is small, the change in the feature representation is minimal.
C. Proof of Theorem 1
According to Lemma 1, the feature extractor cannot distinguish between $\phi(\hat{\bm{x}})$ and $\phi(\bm{x})$, so the model cannot establish an effective decision boundary in the feature space to recognize the trigger pattern.
Next, we analyze the gradient updates of poisoned samples and clean samples separately.
For a poisoned sample $(\hat{\bm{x}},y)$ and a clean sample $(\bm{x},y)$, we compute the gradient of the loss function with respect to the model parameters $\theta$:
$\nabla_{\theta}L(f_{\theta}(\hat{\bm{x}}),y)$ $=\nabla_{\theta}L(f_{\theta}(\bm{x}+\delta),y)$
$=\nabla_{f}L(f_{\theta}(\bm{x}+\delta),y)\cdot\nabla_{\theta}f_{\theta}(\bm{x}+\delta)$
Similarly,
$\nabla_{\theta}L(f_{\theta}(\bm{x}),y)=\nabla_{f}L(f_{\theta}(\bm{x}),y)\cdot\nabla_{\theta}f_{\theta}(\bm{x})$
Since $\delta$ is very small and $f_{\theta}$ is differentiable with respect to $\bm{x}$, we have:
$\nabla_{\theta}f_{\theta}(\bm{x}+\delta)\approx\nabla_{\theta}f_{\theta}(\bm{x})$
and
$f_{\theta}(\bm{x}+\delta)\approx f_{\theta}(\bm{x})+\nabla_{\bm{x}}f_{\theta}(\bm{x})\delta$
Therefore,
$\nabla_{\theta}L(f_{\theta}(\hat{\bm{x}}),y)\approx\nabla_{\theta}L(f_{\theta}(\bm{x}),y).$
Conclusion: The gradient updates from the poisoned samples are almost identical to those from the clean samples with respect to the model parameters, thus they cannot provide additional information to learn the trigger pattern.
D. Static Trigger Patterns and Details
Input: Original trace sequence $\bm{x}$, total number of incoming cells $\Delta_{\text{total}}$, burst size $\delta_{m}$
Output: Poisoned trace $\hat{\bm{x}}$
1
2Initialize the set of candidate insertion locations $\bm{k}_{\text{pool}}$ by randomly sampling $20$ positions from $\bm{x}$;
3 Set initial best insertion locations $\bm{k}=[k_{1},k_{2},\ldots,k_{m}]$;
[…]
5
6for $\text{iteration}=1$ to num_iterations do
7       foreach insertion location $k\in\bm{k}_{\text{pool}}$ do
8             Create modified trace $\hat{\bm{x}}$ by inserting $\delta_{m}$ bursts at $k$;
9             Compute the difference diff between $\bm{x}$ and $\hat{\bm{x}}$ using the statistical distance $D_{F}$;
10
[…]
13                   Update best insertion locations $\bm{k}$;
14
15
16
17return $\hat{\bm{x}}$ with optimized insertion locations $\bm{k}$ and burst sizes $\delta_{m}$;
Algorithm 1 Greedy Optimization for Static Trigger Pattern
The Algorithm 1 employs a greedy optimization strategy to identify the best insertion points for a static trigger pattern within a fully accessible trace sequence $\bm{x}$. Initially, a set of 20 candidate insertion locations, $\bm{k}_{\text{pool}}$, is randomly sampled to ensure a diverse distribution across the trace. The algorithm begins with an empty set of optimal insertion points, $\bm{k}$, and an initial maximum difference, max_diff, set to zero. In each iteration, it evaluates each candidate location $k$ in $\bm{k}_{\text{pool}}$ by generating a modified trace $\hat{\bm{x}}$ with inserted bursts of incoming cells, $\delta_{m}$, at each location. The statistical distance $D_{F}$ between $\bm{x}$ and $\hat{\bm{x}}$ is computed, and if this calculated difference exceeds the current max_diff, the algorithm updates max_diff and sets the best insertion locations to $\bm{k}$. After all iterations, the final output is the optimized poisoned trace $\hat{\bm{x}}$ with insertion locations $\bm{k}$ that maximize distinguishability for effective defense under static data conditions.
E. Dynamic Trigger Patterns and Details
The dynamic trigger pattern is divided into two parts: training and inference.
Input: Training dataset $\bm{x}$, total insertion length $\Delta_{\text{total}}$, model $h$ with LSTM architecture
Output: Trained model $h$ for dynamic trigger prediction
1
2foreach sample $\bm{x}$ in training dataset do
3       Randomly sample a set of insertion locations $\{k_{m}\}$ from $\bm{x}$;
4       foreach insertion location $k_{m}$ do
5             Extract input sequence $\bm{x}[:k_{m}]$ up to insertion point $k_{m}$;
6             Predict burst length $\delta_{m}=h(\bm{x}[:k_{m}])$;
7
8      Generate modified trace $\hat{\bm{x}}$ by inserting predicted bursts $\delta_{m}$ at each $k_{m}$;
9       Compute sequence difference loss:
$L_{\text{sequence}}=-D_{F}(\bm{x},\hat{\bm{x}})$
Compute constraint loss to ensure total insertion constraint $\Delta_{\text{max}}$:
$L_{\text{constraint}}=\left(\text{sum}(\bm{\delta})-\Delta_{\text{max}}\right)^{2}$
Compute total loss:
$L_{\text{total}}=L_{\text{sequence}}+\lambda\cdot L_{\text{constraint}}$
Update model parameters by backpropagating $L_{\text{total}}$;
10
11return Trained model $h$;
Algorithm 2 Dynamic Trigger Pattern Training
The training Algorithm 2 for the dynamic trigger pattern model $h$ involves using random insertion points $\{k_{m}\}$ in each sample trace to simulate dynamic conditions in WF defenses. For each insertion point, the model takes the sequence up to $k_{m}$, $\bm{x}[:k_{m}]$, and predicts an optimal burst length $\delta_{m}$. The model is trained to maximize the sequence difference $L_{\text{sequence}}=-D_{F}(\bm{x},\hat{\bm{x}})$ while adhering to an insertion limit via $L_{\text{constraint}}$, thus equipping it to dynamically adjust trigger patterns in the data dynamic scenario.
In training the dynamic trigger pattern generator, we use the Rimmer dataset with a learning rate of $4\times 10^{-6}$ and a batch size of 1024, running for one epoch on an NVIDIA A40 GPU with Ubuntu 20.04.
Input: Test trace $\bm{x}$, trained model $h$
Output: Poisoned trace $\hat{\bm{x}}$ with inserted triggers
1
2Initialize $\hat{\bm{x}}\leftarrow\bm{x}$;
3 Select insertion points $\{k_{m}\}$ as required;
4 foreach insertion location $k_{m}$ do
5       Extract preceding sequence $\bm{x}[:k_{m}]$ up to $k_{m}$;
6       Predict burst length $\delta_{m}=h(\bm{x}[:k_{m}])$;
7       Insert $\delta_{m}$ dummy packets at location $k_{m}$ in $\hat{\bm{x}}$;
8
9return Poisoned trace $\hat{\bm{x}}$ with dynamic trigger pattern;
Algorithm 3 Dynamic Trigger Pattern Inference
The inference Algorithm 3 operates in scenarios with limited trace access, where the defender cannot pre-optimize insertion points. During inference, a set of random insertion points $k_{m}$ is selected as needed. For each $k_{m}$, the preceding portion of the trace, $\bm{x}[:k_{m}]$, is fed into the dynamic trigger prediction model $h$, which outputs the burst length $\delta_{m}$ for that insertion. The predicted burst lengths are then used to insert the specified packets, thereby modifying the trace in real time. This flexible insertion process supports effective defense under restricted-access conditions.

### v2

@@ VI-G Measuring Adversarial Misclassification Effect
TABLE VI: The overhead and performance of defenses with the clean model under the Sirinam dataset.
Defense Overhead (%) $\downarrow$ Accuracy (%) $\downarrow$
@@
Undefended 0.0 0.0 99.0 98.8 98.3 98.6
TrapFlow 73.7 0.0 4.7 2.8 3.4 5.9
TrapFlow 71.4 0.0 17.1 11.9 15.2 10.8
In this subsection, we evaluate the effectiveness of TrapFlow against four mainstream WF attack models (RF, DF, TikTok, VarCNN) using clean, unpoisoned models, where the trigger is only injected at test time (TrapFlow). The experimental results are shown in Tab. VI. The following conclusions can be drawn from the results in the table: ❶ Even when the attacker’s model is trained without any poisoning, TrapFlow can still significantly degrade the attack performance. For example, the recognition accuracy under the VarCNN attack drops from 99.0% to 10.8%. This confirms the strong adversarial property of our trigger: without any knowledge of the attacker’s model or access to its training process, TrapFlow can still reliably induce domain shifts at test time, severely impairing the attacker’s traffic recognition capability. ❷ Comparing the fully adapted TrapFlow (trained with trigger-injected data) and the clean-model TrapFlow, we observe that although the defense effect is slightly weaker in the clean model setting (e.g., 17.1% vs. 4.7% in RF), TrapFlow still achieves substantial accuracy reduction across all attacks without requiring any training set access.
Why retraining fails. Although an attacker can adapt to the defense mechanism by repeatedly visiting the same website and retraining based on newly collected traffic, this process is in practice equivalent to a backdoor learning process rather than an effective removal or evasion strategy. First, the randomness of the perturbations in terms of their insertion locations makes it difficult to form stable and consistent learnable patterns at the site level, thereby increasing the difficulty of modeling or explicitly removing the defense patterns during retraining.
More critically, these perturbations are not tied to site-specific features but remain stably associated with the target label in a statistical sense, thus exhibiting site-level inconsistency but label-level consistency across visits. Even if the attacker continuously collects new data and retrains, the perturbed samples observed by the model at the website level are still difficult to be internalized as new discriminative features; on the contrary, the label-level bias introduced by the defense is further reinforced during training, such that retraining not only fails to recover the attacking performance, but instead strengthens the association between the perturbations and the target labels. This mechanism is fundamentally different from defense methods that rely on random noise or obfuscation strategies, as it instead exploits the attacker’s own adaptive training process to establish a training-phase defense effect.
VII Discussion and Analysis
@@
Exploration of optimizing trigger patterns for different target labels. In Subsection VI-F, we have explored the potential impact of the choice of target label on the effectiveness of TrapFlow defense. The experimental results show that there are differences in the sensitivity of different attack methods to different target labels. Although we reduce the impact of this difference on TrapFlow defense performance by randomly selecting ten different labels in our experiments, the choice of target labels may introduce minor variations in defense effectiveness, typically within a 2-3% range. We hypothesize that although most label choices do not significantly weaken the defense performance of TrapFlows, in some cases, the wrong choice of target labels may lead to a decrease in the defense effectiveness of TrapFlows. For this reason, it is important for future research to fully understand how the different target labels affect the effectiveness of defenses and to come up with TrapFlow defense strategies that can change based on the labels.
Scalability and deployment limitations in real Tor networks. While TrapFlow supports real-time deployment (Section VI-D) at Tor exit nodes without requiring access to attacker models or training data, it still faces system-level challenges when scaled to large, real-world environments. In practice, Tor exit nodes vary significantly in bandwidth capacity, latency characteristics, and policy constraints. Our current implementation assumes static deployment on selected nodes, which may limit coverage. To achieve broader protection, future deployment may require coordinated trigger injection strategies and adaptive traffic scheduling across multiple exit nodes, while carefully managing bandwidth overhead. Exploring centralized orchestration or decentralized coordination mechanisms could enable more flexible and scalable deployment, enhancing robustness under dynamic network and adversarial conditions.
VIII Conclusion
@@
This paper proposed TrapFlow, a novel Website Fingerprint Defense based on dynamic backdoor learning. TrapFlow embeds backdoor triggers in the attacker’s model to directly control its output, achieving high performance with low overhead. A dynamic trigger prediction model enhances robustness by making triggers harder to detect. Experiments demonstrate that TrapFlow significantly outperforms state-of-the-art defenses, reducing RF’s accuracy from 99% to 6% with 74% data overhead. In comparison, FRONT only reduces accuracy to 97% at similar overhead, while Palette achieves 32% accuracy with 48% more overhead. We further prototyped TrapFlow using WFDefProxy and verified its effectiveness in the real Tor network.